# Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics

### v1

@@
Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics
@@
(Preprint of September 13, 2024)
Extreme mass ratio inspirals (EMRIs) are anticipated to be primary gravitational wave sources for the Laser Interferometer Space Antenna (LISA). They form in dense nuclear clusters, when a compact object is captured by the central massive black holes (MBHs) as a consequence of the frequent two-body interactions occurring between orbiting objects. The physics of this process is complex and requires detailed statistical modeling of a multi-body relativistic system. We present a novel Monte Carlo approach to evolve the post-Newtonian (PN) equations of motion of a compact object orbiting an MBH accounting for the effects of two-body relaxation locally on the fly, without leveraging on the common approximation of orbit-averaging. We apply our method to study the function $S(a_{0})$, describing the fraction of EMRI to total captures (including EMRIs and direct plunges, DPs) as a function of the initial semi-major axis $a_{0}$ for compact objects orbiting central MBHs with $M_{\bullet}\in[10^{4}\,{\rm M}_{\sun},4\times 10^{6}\,{\rm M}_{\sun}]$. The past two decades consolidated a picture in which $S(a_{0})\rightarrow 0$ at large initial semi-major axes, with a sharp transition from EMRI to DPs occurring around a critical scale $a_{\rm c}$. A recent study challenges this notion for low-mass MBHs, finding EMRIs forming at $a\gg a_{\rm c}$ which were named “cliffhangers”. Our simulations confirm the existence of cliffhanger EMRIs, which we find to be more common then previously inferred. Cliffhangers start to appear for $M_{\bullet}\lesssim 3\times 10^{5}\,{\rm M}_{\sun}$ and can account for up to 55% of the overall EMRIs forming at those masses. We find $S(a_{0})\gg 0$ for $a\gg a_{\rm c}$, reaching values as high as 0.6 for $M_{\bullet}=10^{4}\,{\rm M}_{\sun}$, much larger than previously found. We test how these results are influenced by different assumptions on the dynamics used to evolve the system and treatment of two-body relaxation. We find that the PN description of the system greatly enhances the number of EMRIs by shifting $a_{\rm c}$ to larger values at all MBH masses. Conversely, the local treatment of relaxation has a mass dependent impact, significantly boosting the number of cliffhangers at small MBH masses compared to an orbit-averaged treatment. These findings highlight the shortcomings of standard approximations used in the EMRI literature and the importance of carefully modeling the (relativistic) dynamics of these systems. The emerging picture is more complex than previously thought, and should be considered in future estimates of rates and properties of EMRIs detectable by LISA.
@@ 1 Introduction
Most galaxies host a massive black hole (MBH) at their core, with masses in the range of $10^{5}\text{-}10^{10}\,\mathrm{M}_{\sun}$ (Kormendy & Richstone 1995; Magorrian et al. 1998; Kormendy & Ho 2013). Often, these MBHs sit at the centre of nuclear clusters (Carollo et al. 1997; Böker et al. 2002; Côté et al. 2006; Schödel et al. 2007, 2009; Turner et al. 2012): highly dense clusters of stars and compact objects (COs) observed in the nuclei of many galaxies. In the central region of nuclear clusters, densities can reach up to $10^{7}\,\mathrm{M}_{\sun}\mathrm{pc}^{-3}$ (Neumayer et al. 2020), facilitating the scattering of COs onto tight and eccentric orbits around the central MBH, thus forming with it binaries with extreme mass ratios (Binney & Tremaine 2008; Merritt 2013a).
@@
General relativity (GR) predicts that these binary systems emit gravitational waves (GWs), mostly during each CO pericentre passage (Misner et al. 1973). Over time, the CO orbit shrinks and circularises (Peters 1964) as a consequence of the loss of energy and angular momentum of the binary system, which are carried away by GWs (Maggiore 2007). Accordingly, the CO slowly inspirals towards the MBH, completing thousands orbital cycles, eventually merging with it (Amaro-Seoane 2018). These GW sources are known as extreme mass ratio inspirals (EMRIs) due to the large mass imbalance between the pair of objects that make up the binary system: the MBH and the orbiting CO (Amaro-Seoane et al. 2007; Amaro Seoane 2022). Alternatively, if the merger occurs after few orbits or in a head-on collision, the event is referred to as a direct plunge (DP).
@@
EMRIs that form around MBHs with masses in the order of $10^{4}\text{-}10^{7}\,\mathrm{M}_{\sun}$ are expected to emit GWs in the mHz frequency range (Barack & Cutler 2004; Amaro-Seoane 2018). The numerous bursts of GWs emitted during an EMRI are anticipated to build up a significant signal-to-noise ratio (Barack & Cutler 2004; Babak et al. 2017), enabling the detection of such events by the upcoming Laser Interferometer Space Antenna (LISA, Amaro-Seoane et al. 2017, 2023; Colpi et al. 2024). Detecting EMRIs will enable us to test GR (Ryan 1997; Barack & Cutler 2007), estimate cosmological parameters (MacLeod & Hogan 2008; Laghi et al. 2021), and investigate dormant MBHs (Gair et al. 2010; Gallo & Sesana 2019).
@@
Recent works (e.g. Metzger et al. 2022; Franchini et al. 2023; Linial & Metzger 2023) have also proposed that EMRIs might be responsible for the novel, currently unexplained, X-ray variability phenomenon of quasi-periodic eruptions (Miniutti et al. 2019; Giustini et al. 2020; Arcodia et al. 2021, 2024) and recent observations agree with such interpretation (Nicholl et al. 2024). This would offer a way to detect and study EMRIs within the electromagnetic spectrum, enabling multi-messenger observations that would boost their potential as astrophysical and cosmological probes (e.g. Toubiana et al. 2021).
@@
EMRIs can form in many different ways, including: (i) the tidal separation of a stellar-mass black hole (BH) binary system by the central MBH followed by capture of one of the binary components (Hills 1988; Miller et al. 2005); (ii) formation or capture and subsequent migration of BHs within accretion discs surrounding MBHs (Levin 2007; Pan & Yang 2021); (iii) capture of a BH remnant by an MBH as a consequence of supernova kicks (Bortolas et al. 2017; Bortolas & Mapelli 2019); (iv) scatter of BHs onto tight orbits due to the gravitational influence of a secondary MBH in the aftermath of galaxy mergers (Chen et al. 2011; Bode & Wegg 2014; Naoz et al. 2022; Mazzolari et al. 2022).
@@
Still, the standard and ubiquitous formation channel for EMRI formation, which we explore in this work, is the scatter of a CO onto a sufficiently tight and eccentric orbit as a consequence of two-body relaxation (Alexander 2017; Amaro-Seoane 2018). The estimated EMRI formation rate through this channel is in the range of $10^{-8}\text{-}10^{-6}$ EMRIs per year for a central MBH of mass $M_{\bullet}=10^{6}\,\mathrm{M}_{\sun}$ (Hils & Bender 1995; Gair et al. 2004; Rubbo et al. 2006). Considering also the uncertainty in the number density of MBHs inside the observable range of LISA and the fraction of these which are hosted in nuclear clusters, the predicted number of EMRIs which will be detected by LISA during its lifetime spans over three orders of magnitude, from a few to a few thousand events per year (Babak et al. 2017).
@@
Simulating the formation of EMRIs in this standard channel poses considerable challenges. The most accurate description of the whole nuclear cluster would involve a direct N-body computation of the orbits of all stars and COs within it. However, the need to follow these objects for thousands of orbits, and the large mass ratio between them and the central MBH pose an insurmountable challenge for N-body approaches, which show problems even in the simplest Newtonian regime (Amaro-Seoane 2018). Alternative methods, such as solving the steady-state Fokker-Planck equation for the distribution function of the system (e.g. Amaro-Seoane & Preto 2011; Bar-Or & Alexander 2016; Vasiliev 2017; Broggi et al. 2022; Kaur et al. 2024) or employing Monte Carlo codes (e.g. Hénon 1971; Giersz 1998; Joshi et al. 2000; Freitag & Benz 2001; Pattabiraman et al. 2013; Fragione & Sari 2018; Zhang & Amaro Seoane 2024), offer faster approximated solutions but carry inherent limitations and assumptions.
@@
From the aforementioned body of work a picture emerged in which there exists a critical initial semi-major axis $a_{\mathrm{c}}$ for the CO orbit around the MBH determining its eventual fate (Hopman & Alexander 2005; Bar-Or & Alexander 2016; Raveh & Perets 2021). In short, COs starting their inspiral from orbits with $a_{0}<a_{\mathrm{c}}$ would only result in the formation of EMRIs, while only DPs would form when $a_{0}>a_{\mathrm{c}}$, with a sharp transition between the two regimes around $a_{\mathrm{c}}$. Notably, this finding has been derived from detailed simulations of nuclear clusters hosting an MBH with $M_{\bullet}\geq 10^{6}\,{\rm M}_{\odot}$ and then extrapolated at lower masses. Recently, however, such dichotomy has been questioned in Qunbar & Stone (2023) (hereinafter QS23) for the case of low-mass MBHs and intermediate-mass BHs, opening to the possibility of EMRIs forming even from initially wide orbits, dubbed cliffhanger EMRIs. This is an intriguing claim that might bear interesting consequences for estimating LISA EMRI detection rates.
@@
As mentioned above, the techniques and codes employed to verify both the existence of $a_{\mathrm{c}}$ and the appearance of cliffhanger EMRIs rely on a number of assumptions and approximations. A notable one is that of being allowed to average the effects on the orbit of two-body relaxation over an entire orbital period (e.g. Cohn & Kulsrud 1978; Hopman & Alexander 2005; Qunbar & Stone 2023). However, such simplification might not accurately describe the system, as two-body encounters strongly depend on the instantaneous velocity of the orbiting CO (Binney & Tremaine 2008; Merritt 2013a), which significantly changes during a highly eccentric orbit. Another one is to use Newtonian dynamics to describe the system. GR effects are usually included only by adding extra orbit-averaged loss terms due to GW emission and by adjusting the CO capture condition by a suitable factor.
@@
In this work we present a novel post-Newtonian (PN) Monte Carlo code that evolves the PN dynamics of a CO orbiting an MBH. The code treats two-body relaxation locally, by applying multiple small kicks to the CO orbit within a single orbit, and can account for an arbitrary number of stellar components contributing to relaxation and to the potential in which the CO is evolving. As a first application, here we apply this code to study the fraction of EMRIs relative to DPs that form around MBHs of different masses as a function of the initial semi-major axis of the orbit $a_{0}$. This allows us to verify the existence of $a_{\mathrm{c}}$ and its dependence on the system parameters, and the emergence of the novel population of cliffhanger EMRIs found in QS23.
@@
The study presented in QS23 has some limitations which our code can lift. It employs a Monte Carlo approach to evolve in time the Newtonian energy and angular momentum of an orbit subjected to two-body relaxation, which is orbit-averaged over more than a full orbit at times. The loss of energy and angular momentum due to GW radiation is accounted for through equations detailed in Peters (1964), which are orbit-averaged as well. Moreover, only two-body encounters between a stellar-mass BH of mass $m_{\bullet}=10\,\mathrm{M}_{\sun}$ and a single-mass population of $m_{\star}=1\,\mathrm{M}_{\sun}$ stars surrounding the central MBH are considered. The stars are assumed to follow a simple isotropic power-law numeric density, and their global gravitational potential is not considered. In comparison, in this work we track the orbiting of a stellar-mass BH in the total potential well of an MBH and of two populations of objects surrounding it: stars with masses $m_{\star}$ and stellar-mass BHs with masses $m_{\bullet}$. We describe both populations with Dehnen numeric density profiles (Dehnen 1993; Tremaine et al. 1994). We also solve the equations of motion up to the 2.5PN term [^1] In what follows, we call “$n$PN term” the term of the Hamiltonian defined in Eq. (41) which is suppressed by a factor $(1/c)^{2n}$ with respect to the leading one., thus naturally accounting for GW radiation, and locally correct the shape of the orbit for two-body relaxation multiple times within each orbital period.
@@
The paper is organised as follows. In Sect. 2, we review aspects of theory relevant to our implementation and predict some behaviours of cliffhanger EMRIs with analytical means. In Sect. 3, we delve into the specifics of our methods, to then present the results of our simulations in Sect. 4. Finally, in Sect. 5 we discuss our findings and draw our conclusions.
@@ 2 Theory of relaxation-driven EMRI formation
In this section, we first introduce some notions on the treatment of two-body interactions in the loss cone framework. Then, we give operative definitions for EMRIs and DPs. Finally, we briefly review the literature on the study of the EMRI-to-DP ratio and draw some first theoretical results on cliffhanger EMRIs.

### v2

# Hanging on the cliff: Extreme mass ratio inspiral formation with local two-body relaxation and post-Newtonian dynamics

@@
Hanging on the cliff: Extreme mass ratio inspiral formation with local two-body relaxation and post-Newtonian dynamics
@@
(Received 19 September 2024 / Accepted 20 January 2025)
Extreme mass ratio inspirals (EMRIs) are anticipated to be primary gravitational wave sources for the Laser Interferometer Space Antenna (LISA). They form in dense nuclear clusters when a compact object is captured by the central massive black holes (MBHs) as a consequence of the frequent two-body interactions occurring between orbiting objects. The physics of this process is complex and requires detailed statistical modelling of a multi-body relativistic system. We present a novel Monte Carlo approach to evolving the post-Newtonian (PN) equations of motion of a compact object orbiting an MBH. The approach accounts for the effects of two-body relaxation locally on the fly, without leveraging on the common approximation of orbit-averaging. We applied our method to study the function $S(a_{0})$, describing the fraction of EMRI to total captures (including EMRIs and direct plunges, DPs) as a function of the initial semi-major axis $a_{0}$ for compact objects orbiting central MBHs with $M_{\bullet}\in[10^{4}\,{\rm M}_{\sun},4\times 10^{6}\,{\rm M}_{\sun}]$. The past two decades have consolidated a picture in which $S(a_{0})\rightarrow 0$ at large initial semi-major axes, with a sharp transition from EMRIs to DPs occurring around a critical scale $a_{\rm c}$. A recent study challenges this notion for low-mass MBHs, finding EMRIs forming at $a\gg a_{\rm c}$, which were called ‘cliffhangers’. Our simulations confirm the existence of cliffhanger EMRIs, which we find to be more common then previously inferred. Cliffhangers start to appear for $M_{\bullet}\lesssim 3\times 10^{5}\,{\rm M}_{\sun}$ and can account for up to 55% of the overall EMRIs forming at those masses. We find $S(a_{0})\gg 0$ for $a\gg a_{\rm c}$, reaching values as high as 0.6 for $M_{\bullet}=10^{4}\,{\rm M}_{\sun}$, much higher than previously found. We test how these results are influenced by different assumptions on the dynamics used to evolve the system and treatment of two-body relaxation. We find that the PN description of the system greatly enhances the number of EMRIs by shifting $a_{\rm c}$ to larger values at all MBH masses. Conversely, the local treatment of relaxation has a mass-dependent impact, significantly boosting the number of cliffhangers at low MBH masses compared to an orbit-averaged treatment. These findings highlight the shortcomings of standard approximations used in the EMRI literature and the importance of carefully modelling the (relativistic) dynamics of these systems. The emerging picture is more complex than previously thought, and should be considered in future estimates of rates and properties of EMRIs detectable by LISA.
@@ 1 Introduction
Most galaxies host a massive black hole (MBH) at their core, with masses in the range of $10^{5}\text{-}10^{10}\,\mathrm{M}_{\sun}$ (Kormendy & Richstone 1995; Magorrian et al. 1998; Kormendy & Ho 2013). Often these MBHs are found at the centre of nuclear clusters (Carollo et al. 1997; Böker et al. 2002; Côté et al. 2006; Schödel et al. 2007, 2009; Turner et al. 2012): highly dense clusters of stars and compact objects (COs) observed in the nuclei of many galaxies. In the central region of nuclear clusters, densities can reach up to $10^{7}\,\mathrm{M}_{\sun}\mathrm{pc}^{-3}$ (Neumayer et al. 2020), facilitating the scattering of COs onto tight and eccentric orbits around the central MBH, thus forming binaries with extreme mass ratios (Binney & Tremaine 2008; Merritt 2013a).
@@
General relativity (GR) predicts that these binary systems emit gravitational waves (GWs), mostly during each CO pericentre passage (Misner et al. 1973). Over time the CO orbit shrinks and circularises (Peters 1964) as a consequence of the loss of energy and angular momentum of the binary system, which are carried away by GWs (Maggiore 2007). Accordingly, the CO slowly inspirals towards the MBH, completing thousands of orbital cycles, eventually merging with it (Amaro-Seoane 2018). These GW sources are known as extreme mass ratio inspirals (EMRIs) due to the large mass imbalance between the pair of objects that make up the binary system: the MBH and the orbiting CO (Amaro-Seoane et al. 2007; Amaro Seoane 2022). Alternatively, if the merger occurs after few orbits or in a head-on collision, the event is referred to as a direct plunge (DP).
@@
Extreme mass ratio inspirals that form around MBHs with masses of the order of $10^{4}\text{-}10^{7}\,\mathrm{M}_{\sun}$ are expected to emit GWs in the millihertz frequency range (Barack & Cutler 2004; Amaro-Seoane 2018). The numerous bursts of GWs emitted during an EMRI are anticipated to build up a significant signal-to-noise ratio (Barack & Cutler 2004; Babak et al. 2017), enabling the detection of such events by the upcoming Laser Interferometer Space Antenna (LISA, Amaro-Seoane et al. 2017, 2023; Colpi et al. 2024). Detecting EMRIs will enable us to test GR (Ryan 1997; Barack & Cutler 2007; Ghosh & Chakravarti 2024), estimate cosmological parameters (MacLeod & Hogan 2008; Laghi et al. 2021), and investigate dormant MBHs (Gair et al. 2010; Gallo & Sesana 2019).
@@
Recent works (e.g. Metzger et al. 2022; Franchini et al. 2023; Linial & Metzger 2023) have also proposed that EMRIs might be responsible for the novel, currently unexplained, X-ray variability phenomenon of quasi-periodic eruptions (Miniutti et al. 2019; Giustini et al. 2020; Arcodia et al. 2021, 2024), and recent observations agree with such an interpretation (Nicholl et al. 2024). This would offer a way to detect and study EMRIs within the electromagnetic spectrum, enabling multi-messenger observations that would boost their potential as astrophysical and cosmological probes (e.g. Toubiana et al. 2021).
@@
Extreme mass ratio inspirals can form in many different ways, including (i) the tidal separation of a stellar-mass black hole (BH) binary system by the central MBH followed by the capture of one of the binary components (Hills 1988; Miller et al. 2005); (ii) the formation or capture and subsequent migration of BHs within accretion discs surrounding MBHs (Levin 2007; Pan & Yang 2021); (iii) the capture of a BH remnant by an MBH as a consequence of supernova kicks (Bortolas et al. 2017; Bortolas & Mapelli 2019); and (iv) the scatter of BHs onto tight orbits due to the gravitational influence of a secondary MBH in the aftermath of galaxy mergers (Chen et al. 2011; Bode & Wegg 2014; Naoz et al. 2022; Mazzolari et al. 2022).
@@
The standard and ubiquitous formation channel for EMRI formation, which we explore in this work, is the scatter of a CO onto a sufficiently tight and eccentric orbit as a consequence of two-body relaxation (Alexander 2017; Amaro-Seoane 2018). The estimated EMRI formation rate through this channel is in the range of $10^{-8}\text{-}10^{-6}$ EMRIs per year for a central MBH of mass $M_{\bullet}=10^{6}\,\mathrm{M}_{\sun}$ (Hils & Bender 1995; Gair et al. 2004; Rubbo et al. 2006). Considering also the uncertainty in the number density of MBHs inside the observable range of LISA and the fraction of these that are hosted in nuclear clusters, the predicted number of EMRIs that will be detected by LISA during its lifetime spans three orders of magnitude, from a few events per year to a few thousand (Babak et al. 2017).
@@
Simulating the formation of EMRIs in this standard channel poses considerable challenges. The most accurate description of the whole nuclear cluster would involve a direct N-body computation of the orbits of all stars and COs within it. However, the need to follow these objects for thousands of orbits, and the high mass ratio between them and the central MBH pose an insurmountable challenge for N-body approaches, which show problems even in the simplest Newtonian regime (Amaro-Seoane 2018). Alternative methods, such as solving the steady-state Fokker-Planck equation for the distribution function of the system (e.g. Amaro-Seoane & Preto 2011; Bar-Or & Alexander 2016; Vasiliev 2017; Broggi et al. 2022; Kaur et al. 2024) or employing Monte Carlo codes (e.g. Hénon 1971; Giersz 1998; Joshi et al. 2000; Freitag & Benz 2001; Pattabiraman et al. 2013; Fragione & Sari 2018; Zhang & Amaro Seoane 2024), offer faster approximated solutions, but carry inherent limitations and assumptions.
@@
From the aforementioned body of work a picture has emerged in which there exists a critical initial semi-major axis $a_{\mathrm{c}}$ for the CO orbit around the MBH determining its eventual fate (Hopman & Alexander 2005; Bar-Or & Alexander 2016; Raveh & Perets 2021). In short, COs starting their inspiral from orbits with $a_{0}<a_{\mathrm{c}}$ would only result in the formation of EMRIs, while only DPs would form when $a_{0}>a_{\mathrm{c}}$, with a sharp transition between the two regimes around $a_{\mathrm{c}}$. Notably, this finding has been derived from detailed simulations of nuclear clusters hosting an MBH with $M_{\bullet}\geq 10^{6}\,{\rm M}_{\odot}$ and then extrapolated at lower masses. Recently, however, this dichotomy has been questioned in Qunbar & Stone (2024) (hereafter QS24) for the case of low-mass MBHs and intermediate-mass BHs, opening to the possibility of EMRIs forming even from initially wide orbits, dubbed cliffhanger EMRIs. This is an intriguing claim that might bear interesting consequences for estimating LISA EMRI detection rates.
@@
As mentioned above, the techniques and codes employed to verify both the existence of $a_{\mathrm{c}}$ and the appearance of cliffhanger EMRIs rely on a number of assumptions and approximations. A notable approximation is that of being allowed to average the effects on the orbit of two-body relaxation over an entire orbital period (e.g. Cohn & Kulsrud 1978; Hopman & Alexander 2005, QS24). However, this simplification might not accurately describe the system, as two-body encounters strongly depend on the instantaneous velocity of the orbiting CO (Binney & Tremaine 2008; Merritt 2013a), which significantly changes during a highly eccentric orbit. Another is to use Newtonian dynamics to describe the system. GR effects are usually included only by adding extra orbit-averaged loss terms due to GW emission and by adjusting the CO capture condition by a suitable factor.
@@
In this work we present a novel post-Newtonian (PN) Monte Carlo code that evolves the PN dynamics of a CO orbiting an MBH. The code treats two-body relaxation locally, by applying multiple small kicks to the CO orbit within a single orbit, and can account for an arbitrary number of stellar components contributing to relaxation and to the potential in which the CO is evolving. As a first application, here we apply this code to study the fraction of EMRIs relative to DPs that form around MBHs of different masses as a function of the initial semi-major axis of the orbit $a_{0}$. This allows us to verify the existence of $a_{\mathrm{c}}$ and its dependence on the system parameters, and the emergence of the novel population of cliffhanger EMRIs found in QS24.
@@
The study presented in QS24 has some limitations which our code can improve upon. It employs a Monte Carlo approach to evolve in time the Newtonian energy and angular momentum of an orbit subjected to two-body relaxation, which is orbit-averaged over more than a full orbit at times. The loss of energy and angular momentum due to GW radiation is accounted for through equations detailed in Peters (1964), which are orbit-averaged as well. Moreover, only two-body encounters between a stellar-mass BH of mass $m_{\bullet}=10\,\mathrm{M}_{\sun}$ and a single-mass population of $m_{\star}=1\,\mathrm{M}_{\sun}$ stars surrounding the central MBH are considered. The stars are assumed to follow a simple isotropic power-law numeric density, and their global gravitational potential is not considered. In comparison, in this work we track the orbits of a stellar-mass BH in the total potential well of an MBH and of two populations of objects surrounding it: stars with masses $m_{\star}$ and stellar-mass BHs with masses $m_{\bullet}$. We describe both populations with Dehnen numeric density profiles (Dehnen 1993; Tremaine et al. 1994). We also solve the equations of motion up to the 2.5PN term, [^1] In what follows, we call ‘$n$PN term’ the term of the Hamiltonian defined in Eq. (41), which is suppressed by a factor $(1/c)^{2n}$ with respect to the leading term. thus naturally accounting for GW radiation, and we locally correct the shape of the orbit for two-body relaxation multiple times within each orbital period.
@@
The paper is organised as follows. In Sect. 2 we review the aspects of the theory relevant to our implementation and predict some behaviours of cliffhanger EMRIs with analytical means. In Sect. 3 we delve into the specifics of our methods, to then present the results of our simulations in Sect. 4. Finally, in Sect. 5 we discuss our findings and draw our conclusions.
@@ 2 Theory of relaxation-driven EMRI formation
In this section we first introduce some notions on the treatment of two-body interactions in the loss cone framework. Then, we give operative definitions for EMRIs and DPs. Finally, we briefly review the literature on the study of the EMRI-to-DP ratio and draw some first theoretical results on cliffhanger EMRIs.